OpenStudy (anonymous):

How much energy (in joules) is contained in one mole of photons (electromagnetic quanta) of frequency 8.2E+9 Hz

OpenStudy (anonymous):

@Australopithecus

OpenStudy (anonymous):

We have \(E=N_A\times E_{photon}=N_A\times h\nu\). \(N_A\) is a constant, the Avogadro's number, \(6.022\times 10^{23}\). \(h\) is also a constant, Planck's constant, \(6.626\times 10^{-34}(J.s)\). \(\nu\) is the frequency of photon. Now you can easily calculate the answer.